# Hybrid Beamforming Design for Near-Field ISAC with Modular XL-MIMO

### v1

@@ II-B Transmit ISAC Signal
In particular, we consider a coherent time block consisting of $L$ symbols, during which the communication channels and sensing target parameters are assumed to remain invariant. Let $\mathbf{X}\triangleq[\mathbf{x}[1],\mathbf{x}[2],\ldots,\mathbf{x}[L]]\in\mathbb{C}^{N\times L}$ denote the narrowband transmitted ISAC signal, where $\mathbf{x}[l]$ is the transmitted signal at time index $l$. Let $\mathbf{S}=[\mathbf{s}[1],\mathbf{s}[2],\ldots,\mathbf{s}[L]]\in{\mathbb{C}}^{N_{s}\times L}$ denote the information symbol matrix, where $\mathbf{s}[l]$ is the symbol vector transmitted at time index $l$, and $N_{s}$ is the number of data streams. Therefore, the discrete-time transmitted signal at time index $l$ is given by
@@
where $\mathbf{W}_{\text{RF}}\in\mathbb{C}^{N\times N_{\text{RF}}}$ and ${{\mathbf{W}}_{\text{BB}}}\text{=}\left[{{\mathbf{w}}_{\text{BB},1}},\ldots,{{\mathbf{w}}_{\text{BB},{{N}_{s}}}}\right]\in{{\mathbb{C}}^{{{N}_{\text{RF}}}\times{{N}_{s}}}}$ denote the analog and digital beamformers, respectively. Each entry in $\mathbf{S}$ is assumed to be i.i.d. and Gaussian distributed with zero mean and unit variance. The data streams are assumed to be independent of each other, so we have $\mathbb{E}\left\{\mathbf{s}[l]\mathbf{s}^{H}[l]\right\}=\mathbf{I},\forall l$[5]. Then, the covariance matrix of transmitted signal can be given by
@@
Moreover, owing to the hardware limitations of modular XL-MIMO, the RF chains that are connected to one subarray cannot be connected to other subarrays, leading to a partially connected HBF architecture[8]. The analog beamformer ${{\mathbf{W}}_{\text{RF}}}$ holds the block diagonal structure, which is expressed as
@@
where $\mathbf{\tilde{W}}_{\text{RF}}^{k}=\left[{{\mathbf{w}}_{k1}},{{\mathbf{w}}_{k2}},\ldots,{{\mathbf{w}}_{k{{M}_{\text{RF}}}}}\right]\in\mathbb{C}^{M\times M_{\text{RF}}}$ is the analog beamformer of the $k$-th subarray, ${{\mathbf{w}}_{kt}}$, $\forall t=1,2,\cdots,M_{\text{RF}}$ are the values of the phase shifters at the $k$-th subarray, and each element in ${{\mathbf{w}}_{kt}}$ has unit modulus and continuous phase, i.e., $\left|\mathbf{w}_{kt}[m]\right|^{2}=1,\forall m\in\mathcal{M}$. Additionally, the normalized transmit power constraint is given by $\left\|\mathbf{W}_{\text{RF}}\mathbf{W}_{\text{BB}}\right\|_{F}^{2}\leq N_{s}$.
@@ III-A Communication Performance
As the size of the user’s antenna array is considerably smaller than the transmit XL-array at the BS, i.e., $N_{c}\ll N$, the communication channel between the transmit subarrays and the user’s small receive array can be assumed to exhibit the plane wavefront. On the other hand, due to the large aperture of the transmit antenna array, the spherical-wave propagation needs to be considered among the subarrays.
The planar-wave channel matrix between the $k$-th transmit subarray and the user can be written as[18, 8]:
@@
where ${\mu_{p}^{k}}=\left|\mu_{p}^{k}\right|{{e}^{-j\frac{2\pi}{\lambda}D_{p}^{k}}}$ denotes the complex gain of $p$-th path between the reference antennas of the $k$-th transmit subarray and the user’s receive array, $D_{p}^{k}$ represents the distance between the $k$-th reference antenna and the received reference antenna along the $p$-th multipath, with $p=1$ and $p>1$ representing the line-of-sight (LoS) path and the NLoS path, respectively. $\theta_{cp}^{k}$ and $\theta_{tp}^{k}$ represent the AoA and AoD pairs for the $p$-th path between the $k$-th transmit subarray and the user’s receive array. The array steering vector of the $p$-th path for the $k$-th transmit subarray is denoted as $\mathbf{a}_{tp}^{k}\left(\theta_{tp}^{k}\right)={{\left[1,{{e}^{-j\frac{2\pi}{\lambda}d\sin\theta_{tp}^{k}}},\ldots,{{e}^{-j\frac{2\pi}{\lambda}(M-1)d\sin\theta_{tp}^{k}}}\right]}^{T}}$. Similarly, the array steering vector of the $p$-th path for the user’s receive array is denoted as $\mathbf{a}_{cp}^{k}\left(\theta_{cp}^{k}\right)={{\left[1,{{e}^{-j\frac{2\pi}{\lambda}d\sin\theta_{cp}^{k}}},\ldots,{{e}^{-j\frac{2\pi}{\lambda}({{N}_{c}}-1)d\sin\theta_{cp}^{k}}}\right]}^{T}}$.
Therefore, based on HSPM, the communication channel for single-user multi-stream transmission is given by (10), shown at the top of the page.
@@
In the mmWave/sub-THz band, the channels tend to be more sparse due to significant losses caused by large reflection, diffraction, and scattering effects [19, 20]. As paths with insignificant path gains can be disregarded, the number of multipaths in the mmWave/sub-THz band is quite limited. This limitation can decrease the rank of the communication channel, leading to a decline in the communication spectral efficiency. By making the assumption $KN_{p}\leq N_{c}<N$, we conduct an analysis of the rank of ${{\mathbf{H}}_{c}}$ in Lemma 1 below.
@@
$\mathrm{rank}\left({{\mathbf{H}}_{c}}\right)=KN_{p}.$ (11)
@@
Firstly, for a given propagation path $p$, the subarray response vectors $\mathbf{a}_{tp}^{1}\left(\theta_{tp}^{1}\right),\ldots,\mathbf{a}_{tp}^{K}\left(\theta_{tp}^{K}\right)$ are linearly independent[16]. Furthermore, due to the distinguishability of propagation paths and their varying angles, the response vectors associated with different propagation paths $\mathbf{a}_{t1}^{k}\left(\theta_{t1}^{k}\right),\ldots,\mathbf{a}_{tN_{p}}^{k}(\theta_{tN_{p}}^{k})$ are linearly independent for the $k$-th subarray. Therefore, we can obtain that each row of the communication channel $\mathbf{H}_{c}$ in (10) is a linear combination of $KN_{p}$ linearly independent vectors as
$\mathbf{\bar{a}}_{tp}^{k}{{\left(\phi_{tp}^{k}\right)}^{T}}\triangleq\left[\underbrace{{0},\ldots,{0}}_{k-1},\mathbf{a}_{tp}^{k}{{\left(\phi_{tp}^{k}\right)}^{T}},\underbrace{{0},\ldots,{0}}_{K-k}\right]^{T},$ (12)
where $k=1,2,\cdots,K$, $p=1,2,\cdots,N_{p}$. ∎
This demonstrates that compared to the traditional far-field PWM channel whose rank is upper-bounded by the number of multipaths $N_{p}$[21], the subarray-based near-field HSPM channel possesses a more sufficient rank, thus enabling it to provide more spatial multiplexing gains.
@@
where ${{\mathbf{Z}}_{c}}\in\mathbb{C}^{N_{c}\times L}$ is a zero-mean complex Gaussian noise matrix with covariance matrix $\sigma_{c}^{2}\mathbf{I}_{N_{c}}$. Then, the achievable communication spectral efficiency can be calculated as[22]
@@ III-B Sensing Performance
For radar sensing, we assume that the BS detects the target of interest located at $\mathbf{l}_{1}=(r_{1}\sin{\phi}_{1},r_{1}\cos{\phi}_{1})$. Assuming there exist $Q-1$ signal-dependent uncorrelated interferences located at $\mathbf{l}_{q}=(r_{q}\sin{\phi}_{q},r_{q}\cos{\phi}_{q}),\forall q\in\{2,3,\cdots,Q\}$, the received signal at the BS over $L$ symbols can be written as
@@
${{\mathbf{Y}}_{s}}=\underbrace{{{\beta}_{{1}}}{{\mathbf{g}}_{r1}}\mathbf{g}_{t1}^{H}\mathbf{X}}_{\text{Target reflection}}+\underbrace{\sum\nolimits_{q=2}^{Q}{{{\beta}_{q}}{{\mathbf{g}}_{rq}}\mathbf{g}_{tq}^{H}\mathbf{X}}}_{\text{Echosignal of interferences}}+{{\mathbf{Z}}_{s}}$ (15)
@@
where $\beta_{q}$ is the complex reflection coefficient proportional to the radar cross section (RCS) of the $q$-th object with $\mathbb{E}\left\{\left|\beta_{q}\right|^{2}\right\}=\alpha_{q}^{2}$ ($q=1$ for the target and $q\neq 1$ for the interferences), and ${{\mathbf{Z}}_{s}}=\left[{{\mathbf{z}}_{1}},{{\mathbf{z}}_{2}},\ldots,{{\mathbf{z}}_{L}}\right]\in{{\mathbb{C}}^{N\times L}}$ denotes the additive white Gaussian noise (AWGN), with each column being independent and i.i.d. circularly symmetric complex Gaussian random vector with zero mean and covariance $\mathbf{R}_{s}=\sigma_{s}^{2}\mathbf{I}_{N}$. Additionally, ${{\mathbf{g}}_{rq}}$ and ${{\mathbf{g}}_{tq}}$ denote the receive and transmit array response vectors for the $q$-th object at BS based on HSPM, respectively, which can be expressed as
${{\mathbf{g}}_{tq}}={{\mathbf{g}}_{t}}\left({{\mathbf{l}}_{q}}\right)=\left(\mathrm{diag}\left({{\bm{\nu}}_{tq}}\right)\otimes{{\mathbf{I}}_{M}}\right){{{\mathbf{\hat{a}}}}_{tq}},$ (16)
${{\mathbf{g}}_{rq}}={{\mathbf{g}}_{r}}\left({{\mathbf{l}}_{q}}\right)=\left(\mathrm{diag}\left({\bm{\nu}_{rq}}\right)\otimes{{\mathbf{I}}_{M}}\right){{{\mathbf{\hat{a}}}}_{rq}},\forall q\in\{1,\cdots,Q\},$
@@
where ${{\mathbf{\hat{a}}}_{tq}}={{[{{(\mathbf{a}_{tq}^{1})}^{T}},{{(\mathbf{a}_{tq}^{2})}^{T}},\ldots,{{(\mathbf{a}_{tq}^{K})}^{T}}]}^{T}}$ and ${{\mathbf{\hat{a}}}_{rq}}={{[{{(\mathbf{a}_{rq}^{1})}^{T}},{{(\mathbf{a}_{rq}^{2})}^{T}},\ldots,{{(\mathbf{a}_{rq}^{K})}^{T}}]}^{T}}$, with $\mathbf{a}_{tq}^{k}={{[1,{{e}^{-j\frac{2\pi}{\lambda}d\sin\phi_{tq}^{k}}},\ldots,{{e}^{-j\frac{2\pi}{\lambda}(M-1)d\sin\phi_{tq}^{k}}}]}^{T}}$ and $\mathbf{a}_{rq}^{k}={{[1,{{e}^{-j\frac{2\pi}{\lambda}d\sin\phi_{rq}^{k}}},\ldots,{{e}^{-j\frac{2\pi}{\lambda}(M-1)d\sin\phi_{rq}^{k}}}]}^{T}}$ denoting the intra-subarray response vectors for the Tx and Rx under the uniform planar-wave assumption, respectively. The angles of $\mathbf{l}_{q}$ as observed from the $k$-th subarray at Tx and Rx are denoted as ${\phi}_{tq}^{k}$ and ${\phi}_{rq}^{k}$, respectively, which can be obtained as
$\sin\phi_{tq}^{k}=\frac{r_{q}\sin\phi_{q}-{{D}_{0}}-\left(k-1\right){{d}_{s}}}{\left\|{{\mathbf{l}}_{q}}-\mathbf{l}_{k}^{t}\right\|},$ (17)
$\sin\phi_{rq}^{k}=\frac{r_{q}\sin\phi_{q}+{{D}_{0}}+\left(k-1\right){{d}_{s}}}{\left\|{{\mathbf{l}}_{q}}-\mathbf{l}_{k}^{r}\right\|}.$
@@
Additionally, ${{\bm{\nu}}_{tq}}$ and ${{\bm{\nu}}_{rq}}$ represent the inter-subarray response vectors of Tx and Rx toward the $q$-th object, respectively, and can be expressed as
@@
${\bm{\nu}}_{rq}=\left[{{e}^{-j\frac{2\pi}{\lambda}\left\|{{\mathbf{l}}_{q}}-\mathbf{l}_{1}^{r}\right\|}},\ldots,{{e}^{-j\frac{2\pi}{\lambda}\left\|{{\mathbf{l}}_{q}}-\mathbf{l}_{K}^{r}\right\|}}\right]^{T}.$
Then, the received signal for the probing target is filtered by the receive beamformer $\mathbf{w}\in\mathbb{C}^{N\times 1}$, and the output of the BS receiver is given as
${\mathbf{y}_{s}}={{\mathbf{w}}^{H}}{{\beta}_{1}}{{\mathbf{g}}_{r1}}\mathbf{g}_{t1}^{H}\mathbf{X}+{{\mathbf{w}}^{H}}\sum\nolimits_{q=2}^{Q}{{{\beta}_{q}}{{\mathbf{g}}_{rq}}\mathbf{g}_{tq}^{H}\mathbf{X}}+{{\mathbf{w}}^{H}}{{\mathbf{Z}}_{s}}.$ (19)
Subsequently, the radar SCNR can be calculated as [23]
${\gamma}_{s}$ $=\frac{E\left[{{\left\|{{\mathbf{w}}^{H}}{{\beta}_{\text{1}}}{{\mathbf{g}}_{r1}}\mathbf{g}_{t1}^{H}\mathbf{X}\right\|}^{2}}\right]}{E\left[{{\left\|{{\mathbf{w}}^{H}}\sum\nolimits_{q=2}^{Q}{{{\beta}_{q}}{{\mathbf{g}}_{rq}}\mathbf{g}_{tq}^{H}\mathbf{X}}\right\|}^{2}}\right]+E\left[{{\left\|{{\mathbf{w}}^{H}}{{\mathbf{Z}}_{s}}\right\|}^{2}}\right]}$ (20)
$=\frac{\alpha_{1}^{2}{{\mathbf{w}}^{H}}{{\mathbf{g}}_{r1}}\mathbf{g}_{t1}^{H}{{\mathbf{R}}_{X}}{{\mathbf{g}}_{t1}}\mathbf{g}_{r1}^{H}\mathbf{w}}{{{\mathbf{w}}^{H}}\left(\sum\nolimits_{q=2}^{Q}{\alpha_{q}^{2}{{\mathbf{g}}_{rq}}\mathbf{g}_{tq}^{H}{{\mathbf{R}}_{X}}{{\mathbf{g}}_{tq}}\mathbf{g}_{rq}^{H}}+\sigma_{s}^{2}{{\mathbf{I}}_{N}}\right)\mathbf{w}}.$
When $\mathbf{R}_{X}$ is given, the optimal $\mathbf{w}^{*}$ to maximize the SCNR can be derived by solving the equivalent minimum variance distortionless response (MVDR) problem[24], where the closed-form optimal solution can be obtained as
${\mathbf{w}^{*}}=\frac{{{\left(\bm{\Sigma}+\sigma_{s}^{2}{{\mathbf{I}}_{N}}\right)}^{-1}}{{\mathbf{g}}_{r1}}}{\mathbf{g}_{r1}^{H}{{\left(\bm{\Sigma}+\sigma_{s}^{2}{{\mathbf{I}}_{N}}\right)}^{-1}}{{\mathbf{g}}_{r1}}},$ (21)
where $\bm{\Sigma}\triangleq\sum\nolimits_{q=2}^{Q}{\alpha_{q}^{2}{{\mathbf{g}}_{rq}}\mathbf{g}_{tq}^{H}{{\mathbf{R}}_{X}}{{\mathbf{g}}_{tq}}\mathbf{g}_{rq}^{H}}$. It is worth noting that once the optimal transmit covariance matrix $\mathbf{R}_{X}$ is derived, the optimal receive beamformer ${\mathbf{w}^{*}}$ can be readily obtained as a closed-form solution from (21). Therefore, we focus on the optimization of transmit beamforming while adopting a fixed receive beamformer $\mathbf{w}$ to avoid the more complex joint transmit-receive beamforming design.

### v2

@@ II-A Transmit ISAC Signal
We consider a coherent time block of $L$ symbols, during which the communication channels and sensing target parameters remain invariant. The narrowband transmitted ISAC signal is denoted by $\mathbf{X}\triangleq[\mathbf{x}[1],\mathbf{x}[2],\ldots,\mathbf{x}[L]]\in\mathbb{C}^{N\times L}$, where $\mathbf{x}[l]$ is the transmitted signal at time index $l$. The information symbol matrix is given by $\mathbf{S}=[\mathbf{s}[1],\mathbf{s}[2],\ldots,\mathbf{s}[L]]\in{\mathbb{C}}^{N_{s}\times L}$, where $\mathbf{s}[l]$ is the symbol vector transmitted at time index $l$, and $N_{s}$ is the number of data streams. Therefore, the discrete-time transmitted signal at time index $l$ is expressed as
@@
where $\mathbf{W}_{\text{RF}}\in\mathbb{C}^{N\times N_{\text{RF}}}$ and ${{\mathbf{W}}_{\text{BB}}}\text{=}\left[{{\mathbf{w}}_{\text{BB},1}},\ldots,{{\mathbf{w}}_{\text{BB},{{N}_{s}}}}\right]\in{{\mathbb{C}}^{{{N}_{\text{RF}}}\times{{N}_{s}}}}$ denote the analog and digital beamformers, respectively. Each entry in $\mathbf{S}$ is assumed to be i.i.d. and Gaussian distributed with zero mean and unit variance. The data streams are assumed to be independent of each other, satisfying $\mathbb{E}\left\{\mathbf{s}[l]\mathbf{s}^{H}[l]\right\}=\mathbf{I},\forall l$[8]. Then, the covariance matrix of transmitted signal is given by
@@
Figure 2: Illustration of the hybrid beamforming architecture for the proposed modular XL-MIMO system.
The analog beamformer ${{\mathbf{W}}_{\text{RF}}}$ exhibits a block diagonal structure:
@@
where $\mathbf{\tilde{W}}_{\text{RF}}^{k}=\left[{{\mathbf{w}}_{k1}},{{\mathbf{w}}_{k2}},\ldots,{{\mathbf{w}}_{k{{M}_{\text{RF}}}}}\right]\in\mathbb{C}^{M\times M_{\text{RF}}}$ is the analog beamformer of the $k$-th subarray, ${{\mathbf{w}}_{kt}}$, $\forall t=1,2,\cdots,M_{\text{RF}}$ are the phase shifter values at the $k$-th subarray, with each element in ${{\mathbf{w}}_{kt}}$ has unit modulus and continuous phase, i.e., $\left|\mathbf{w}_{kt}[m]\right|^{2}=1,\forall m\in\mathcal{M}$. Additionally, the normalized transmit power constraint is given by $\left\|\mathbf{W}_{\text{RF}}\mathbf{W}_{\text{BB}}\right\|_{F}^{2}\leq N_{s}$.
@@ III Performance Metrics and Problem Formulation
This section first introduces the piecewise-far-field channel model for characterizing the communication and sensing channels. Based on this model, we then present the performance metrics for communication and sensing, and formulate the joint analog-digital beamforming optimization problem.
The small subarray aperture resulting from $M\ll N$ limits the near-field range of each subarray to only a few meters, making it more likely for the user, target, or scatterer to be located in the far-field region of each subarray while remaining in the near-field region of the overall array. The small subarray aperture resulting from $M\ll N$ leads to a negligible near-field range for each subarray. Therefore, we adopt the piecewise-far-field model, which assumes far-field propagation within each subarray and near-field propagation among subarrays [19, 10].
Remark 1
For example, considering a modular array with $K=6$, $M=32$ at 38 GHz, the subarray Rayleigh distance is only 2.5 m. In contrast, the overall array has a 40 m Rayleigh distance when collocated (half-wavelength spacing) and 250 m when widely spaced (10-wavelength spacing). Thus, users and targets are typically in the far-field of each subarray but near-field of the overall array.
@@ III-A Communication Performance
As the size of the user’s antenna array is also considerably small, the communication channel between the transmit subarrays and the user’s receive array can be modeled under the far-field assumption. Therefore, the far-field communication channel between the $k$-th transmit subarray and the user is given by[23, 10]:
@@
where $N_{p}$ denotes the total number of distinguishable paths between Tx and the user, including both line-of-sight (LoS) and non-line-of-sight (NLoS) propagation, ${\mu_{p}^{k}}=\left|\mu_{p}^{k}\right|{{e}^{-j\frac{2\pi}{\lambda}D_{p}^{k}}}$ denotes the complex gain of $p$-th path between the $k$-th transmit subarray and the user’s receive array, which captures the near-field effect across different subarrays as it depends on the exact distance $D_{p}^{k}$ between the $k$-th transmit subarray’s reference antenna and the user’s receive array’s reference antenna along the $p$-th multipath, with $p=1$ and $p>1$ representing the LoS path and the NLoS path, respectively. $\theta*{cp}^{k}$ and $\theta*{tp}^{k}$ represent the angle of arrival (AoA) relative to the receive array and angle of departure (AoD) relative to the $k$-th transmit subarray along the $p$-th propagation path between the Tx and user, as illustrated in Fig. 1. The array steering vector of the $p$-th path for the $k$-th transmit subarray is given by $\mathbf{a}_{tp}^{k}\left(\theta_{tp}^{k}\right)={{\left[1,{{e}^{-j\frac{2\pi}{\lambda}d\sin\theta_{tp}^{k}}},\ldots,{{e}^{-j\frac{2\pi}{\lambda}(M-1)d\sin\theta_{tp}^{k}}}\right]}^{T}}$. Similarly, the array steering vector of the $p$-th path for the user’s receive array is given by $\mathbf{a}_{cp}^{k}\left(\theta_{cp}^{k}\right)={{\left[1,{{e}^{-j\frac{2\pi}{\lambda}d\sin\theta_{cp}^{k}}},\ldots,{{e}^{-j\frac{2\pi}{\lambda}({{N}_{c}}-1)d\sin\theta_{cp}^{k}}}\right]}^{T}}$.
Based on the piecewise-far-field channel model, we can characterize the overall communication channel between the user and the transmit array by concatenating the individual far-field subarray channels in (4). The complete mathematical expression is presented in (5) at the bottom of this page.
@@
In the mmWave/sub-THz band, the channels tend to be more sparse due to significant losses caused by large reflection, diffraction, and scattering effects [24, 25]. As paths with insignificant path gains can be disregarded, the number of multipaths in the mmWave/sub-THz band is quite limited. This limitation can decrease the rank of the communication channel, leading to a decline in the communication spectral efficiency. To thoroughly understand this rank-dependent behavior, we provide a rigorous analysis of the rank of ${{\mathbf{H}}_{c}}$ in Lemma 1 below.
@@
$\min\{N_{p},N_{c}\}\leq\mathrm{rank}\left({{\mathbf{H}}_{c}}\right)\leq\min\{KN_{p},N_{c}\}.$ (6)
@@
Please see Appendix A. ∎
Remark 2
For sufficiently large $N_{c}$, the rank of traditional far-field channel is limited by the number of multipaths $N_{p}$ [26]. In contrast, Lemma 1 demonstrates that the subarray-based piecewise-far-field channel model can provide a more sufficient rank that is lower-bounded by $N_{p}$ and can potentially reach $KN_{p}$, enabling the support of up to $KN_{p}$ independent data streams.
@@
where ${{\mathbf{Z}}_{c}}\in\mathbb{C}^{N_{c}\times L}$ is a zero-mean complex Gaussian noise matrix with covariance matrix $\sigma_{c}^{2}\mathbf{I}_{N_{c}}$. Then, the achievable communication spectral efficiency can be calculated as
@@ III-B Sensing Performance
For radar sensing, we consider a target of interest located at $\mathbf{l}_{0}=(r_{0}\sin{\phi}_{0},r_{0}\cos{\phi}_{0})$ and $Q-1$ signal-dependent uncorrelated interferences located at $\mathbf{l}_{q}=(r_{q}\sin{\phi}_{q},r_{q}\cos{\phi}_{q}),\forall q\in\{1,2,\cdots,Q-1\}$. Therefore, the received signal at the BS over $L$ symbols can be expressed as
@@
${{\mathbf{Y}}_{s}}=\underbrace{{{\beta}_{{0}}}{{\mathbf{g}}_{r0}}\mathbf{g}_{t0}^{H}\mathbf{X}}_{\text{Target reflection}}+\underbrace{\sum\nolimits_{q=1}^{Q-1}{{{\beta}_{q}}{{\mathbf{g}}_{rq}}\mathbf{g}_{tq}^{H}\mathbf{X}}}_{\text{Echo signal of interferences}}+{{\mathbf{Z}}_{s}}$ (9)
@@
where $\beta_{q}$ denotes the complex reflection coefficient proportional to the radar cross section (RCS) of the $q$-th object, with $\mathbb{E}\left\{\left|\beta_{q}\right|^{2}\right\}=\alpha_{q}^{2}$ ($q=0$ for the target and $q\geq 1$ for the interferences), the additive white Gaussian noise (AWGN) is represented by ${{\mathbf{Z}}_{s}}=\left[{{\mathbf{z}}_{1}},{{\mathbf{z}}_{2}},\ldots,{{\mathbf{z}}_{L}}\right]\in{{\mathbb{C}}^{N\times L}}$, where each column is an independent and identically distributed (i.i.d.) circularly symmetric complex Gaussian random vector with zero mean and covariance $\mathbf{R}_{s}=\sigma_{s}^{2}\mathbf{I}_{N}$. Moreover, based on the piecewise-far-field channel model, the Rx and Tx array response vectors for the $q$-th object at the BS are denoted by ${{\mathbf{g}}_{rq}}$ and ${{\mathbf{g}}_{tq}}$, respectively, and can be expressed as
${{\mathbf{g}}_{tq}}=\left(\mathrm{diag}\left({{\bm{\nu}}_{tq}}\right)\otimes{{\mathbf{I}}_{M}}\right){{{\mathbf{\hat{a}}}}_{tq}},$ (10a)
${{\mathbf{g}}_{rq}}=\left(\mathrm{diag}\left({\bm{\nu}_{rq}}\right)\otimes{{\mathbf{I}}_{M}}\right){{{\mathbf{\hat{a}}}}_{rq}},$ (10b)
where
${{\mathbf{\hat{a}}}_{tq}}\!=\!\text{vec}\!\left(\left[\mathbf{a}_{tq}^{1},\mathbf{a}_{tq}^{2},\cdots,\mathbf{a}_{tq}^{K}\right]\right),$ (11)
${{\mathbf{\hat{a}}}_{rq}}\!=\!\text{vec}\!\left(\left[\mathbf{a}_{rq}^{1},\mathbf{a}_{rq}^{2},\cdots,\mathbf{a}_{rq}^{K}\right]\right),$
@@
where $\mathbf{a}_{tq}^{k}={{[1,{{e}^{-j\frac{2\pi}{\lambda}d\sin\phi_{tq}^{k}}},\ldots,{{e}^{-j\frac{2\pi}{\lambda}(M-1)d\sin\phi_{tq}^{k}}}]}^{T}}$ and $\mathbf{a}_{rq}^{k}={{[1,{{e}^{-j\frac{2\pi}{\lambda}d\sin\phi_{rq}^{k}}},\ldots,{{e}^{-j\frac{2\pi}{\lambda}(M-1)d\sin\phi_{rq}^{k}}}]}^{T}}$ denote the intra-subarray response vectors for the Tx and Rx, considering the far-field effect within individual subarrays. The angles ${\phi}_{tq}^{k}$ and ${\phi}_{rq}^{k}$ denote the direction of the $q$-th object with respect to the positive $y$-axis, as observed from the perspective of the $k$-th subarray at the Tx and Rx, respectively, which are given by
$\phi_{tq}^{k}=\text{arc}\sin\left(\frac{r_{q}\sin\phi_{q}-{{D}_{0}}-\left(k-1\right){{d}_{s}}}{\left\|{{\mathbf{l}}_{q}}-\mathbf{l}_{k}^{t}\right\|}\right),$ (12a)
$\phi_{rq}^{k}=\text{arc}\sin\left(\frac{r_{q}\sin\phi_{q}+{{D}_{0}}+\left(k-1\right){{d}_{s}}}{\left\|{{\mathbf{l}}_{q}}-\mathbf{l}_{k}^{r}\right\|}\right).$ (12b)
@@
Additionally, ${{\bm{\nu}}_{tq}}$ and ${{\bm{\nu}}_{rq}}$ denote the inter-subarray response vectors of Tx and Rx toward the $q$-th object, characterizing the near-field effect between subarrays, which can be expressed as
@@
${\bm{\nu}}_{rq}=\left[{{e}^{-j\frac{2\pi}{\lambda}\left\|{{\mathbf{l}}_{q}}-\mathbf{l}_{1}^{r}\right\|}},\ldots,{{e}^{-j\frac{2\pi}{\lambda}\left\|{{\mathbf{l}}_{q}}-\mathbf{l}_{K}^{r}\right\|}}\right]^{T},$
where $\mathbf{l}_{k}^{t}$ and $\mathbf{l}_{k}^{r}$ denote the positions of the reference antennas in the $k$-th subarray at Tx and Rx, i.e., $\mathbf{l}_{k}^{t}=\mathbf{l}_{k,1}^{t}$ and $\mathbf{l}_{k}^{r}=\mathbf{l}_{k,1}^{r}$.
Then, the received signal is processed by applying the receive beamformer $\mathbf{w}\in\mathbb{C}^{N\times 1}$, resulting in the following output at the Rx:
${\mathbf{y}_{s}}={{\mathbf{w}}^{H}}{{\beta}_{0}}{{\mathbf{g}}_{r0}}\mathbf{g}_{t0}^{H}\mathbf{X}+{{\mathbf{w}}^{H}}\sum\limits_{q=1}^{Q-1}{{{\beta}_{q}}{{\mathbf{g}}_{rq}}\mathbf{g}_{tq}^{H}\mathbf{X}}+{{\mathbf{w}}^{H}}{{\mathbf{Z}}_{s}}.$ (14)
Subsequently, the sensing SCNR can be calculated as
${\gamma}_{s}=\frac{E\left[{{\left\|{{\mathbf{w}}^{H}}{{\beta}_{0}}{{\mathbf{g}}_{r0}}\mathbf{g}_{t0}^{H}\mathbf{X}\right\|}^{2}}\right]}{E\left[{{\left\|{{\mathbf{w}}^{H}}\sum\nolimits_{q=1}^{Q-1}{{{\beta}_{q}}{{\mathbf{g}}_{rq}}\mathbf{g}_{tq}^{H}\mathbf{X}}\right\|}^{2}}\right]+E\left[{{\left\|{{\mathbf{w}}^{H}}{{\mathbf{Z}}_{s}}\right\|}^{2}}\right]}$ (15)
$=\frac{{{\mathbf{w}}^{H}}{{\mathbf{G}}_{0}}\left({{\mathbf{W}}_{\text{RF}}}{{\mathbf{W}}_{\text{BB}}}\mathbf{W}_{\text{BB}}^{H}\mathbf{W}_{\text{RF}}^{H}\right)\mathbf{G}_{0}^{H}\mathbf{w}}{{{\mathbf{w}}^{H}}\left(\sum\nolimits_{q=1}^{Q-1}{{{\mathbf{G}}_{q}}\left({{\mathbf{W}}_{\text{RF}}}{{\mathbf{W}}_{\text{BB}}}\mathbf{W}_{\text{BB}}^{H}\mathbf{W}_{\text{RF}}^{H}\right)\mathbf{G}_{q}^{H}}+\sigma_{s}^{2}{{\mathbf{I}}_{N}}\right)\mathbf{w}},$
where ${{\mathbf{G}}_{q}}\triangleq\alpha_{q}{{\mathbf{g}}_{rq}}\mathbf{g}_{tq}^{H}$, $q=0,1,\ldots,Q-1$.
When the transmit covariance matrix $\mathbf{R}_{X}$ is given, the optimal receive beamformer $\mathbf{w}^{*}$ that maximizes the SCNR can be obtained by solving the equivalent minimum variance distortionless response (MVDR) problem[27]. The closed-form optimal solution is given by
${\mathbf{w}^{*}}=\frac{{{\left(\bm{\Sigma}+\sigma_{s}^{2}{{\mathbf{I}}_{N}}\right)}^{-1}}{{\mathbf{g}}_{r0}}}{\mathbf{g}_{r0}^{H}{{\left(\bm{\Sigma}+\sigma_{s}^{2}{{\mathbf{I}}_{N}}\right)}^{-1}}{{\mathbf{g}}_{r0}}},$ (16)
where $\bm{\Sigma}\triangleq\sum\nolimits_{q=1}^{Q-1}{\alpha_{q}^{2}{{\mathbf{g}}_{rq}}\mathbf{g}_{tq}^{H}{{\mathbf{R}}_{X}}{{\mathbf{g}}_{tq}}\mathbf{g}_{rq}^{H}}$. It is worth noting that once the optimal transmit covariance matrix $\mathbf{R}_{X}$ is determined, the optimal receive beamformer ${\mathbf{w}^{*}}$ can be readily computed using the closed-form expression in (16). Therefore, in this work, we concentrate on optimizing the transmit beamforming while employing a fixed receive beamformer $\mathbf{w}$ to avoid the more complex joint transmit-receive beamforming design.